I had a question similar to this one earlier. This one is of a different format though. If Given A5= -5 and A9=10 find d and A1
Answer by stanbon(75887)

Given A5= -5 and A9=10 find d and A1
-------------
a(5) = a(1) + 4d = -5
a(9) = a(1) + 8d = 10
------------------------
Subtract 1st from 2nd to get:
4d = 15
d = 15/4
---------
Substitute into a(1)+4d = -5 to solve for a(1)
a(1) +4(15/4) = -5
a(1) + 15 = -5
a(1) = -20
================
Cheers,
Stan H.